### v2

# More Expressive Attention with Negative Weights

Abstract
We propose a novel attention mechanism, named Cog Attention, that enables attention weights to be negative for enhanced expressiveness, which stems from two key factors: (1) Cog Attention can shift the token deletion and copying function from a static OV matrix to dynamic QK inner products, with the OV matrix now focusing more on refinement or modification. The attention head can simultaneously delete, copy, or retain tokens by assigning them negative, positive, or minimal attention weights, respectively. As a result, a single attention head becomes more flexible and expressive. (2) Cog Attention improves the model’s robustness against representational collapse, which can occur when earlier tokens are over-squashed into later positions, leading to homogeneous representations. Negative weights reduce effective information paths from earlier to later tokens, helping to mitigate this issue. We develop Transformer-like models which use Cog Attention as attention modules, including decoder-only models for language modeling and U-ViT diffusion models for image generation. Experiments show that models using Cog Attention exhibit superior performance compared to those employing traditional softmax attention modules. Our approach suggests a promising research direction for rethinking and breaking the entrenched constraints of traditional softmax attention, such as the requirement for non-negative weights. Our code is available at https://github.com/trestad/CogAttn.
Attention with Negative Weights
1 Introduction
The Transformer architecture (Vaswani et al., 2017) has achieved success across numerous applications, such as language modeling (Brown et al., 2020) and image generation (Dosovitskiy et al., 2021). A crucial factor contributing to its success is the softmax attention mechanism (Bahdanau et al., 2015).
Figure 1: In the Indirect Object Identification (IOI) task (Wang et al., 2023), a language model should identify the indirect object ($IO$) from a context that includes both $IO$ and a subject ($S$). Subfigures (a) and (b) show how Cog Attention and softmax attention process a single sequence through a process of elimination, but in different ways: a softmax attention head with a deletion-function OV matrix eliminates all attended tokens. While the $IO$ token receives less attention than $S$, it is also deleted. In contrast, Cog Attention shifts functions like deletion or copying from a static OV matrix to dynamic query-key inner products, allowing the head to assign negative weights to $S$ tokens for elimination while preserving $IO$s. Subfigures (c) and (d) show the attention of these two heads on names across the entire dataset vs the direction of their output embeddings. Cog Attention exhibits a weaker correlation between the attention weights assigned to $IO$s and the extent of their deletion. For further details, please see Section 3.
Softmax ensures non-negative attention weights, but we argue that it limits the expressiveness of the attention mechanism. Figure 1 shows one possible way that negative attention weights enhance the model’s expressiveness using the same number of parameters: in a softmax attention head, the QK matrix (Elhage et al., 2021) determines the relevant tokens for attention, while the OV matrix governs the processing of these attended tokens (e.g., deletion or copying). Suppose a softmax attention head has an OV matrix capable of deleting tokens that the QK matrix attends to; since attention weights must be non-positive, a useful token is also somewhat deleted. By allowing negative attention weights, however, deletion or copying can be expressed through the sign of the attention weight and accomplished during the weighted summation of value vectors. This functional shift also allows the OV matrix to focus more on higher-level tasks, such as refinement or modification, rather than solely handling contextual deletions or copies. Consequently, negative attention weights eliminate irrelevant tokens while preserving useful ones, mitigating the risk of “friendly fire” on useful tokens.
Despite the potential benefits of incorporating negative weights in attention mechanisms, this question has been rarely explored. Apart from the common belief that attention weights should naturally be non-negative, introducing negative weights can lead to challenges such as training instability, numerical overflow, and difficulties in attention normalization due to issues like division by zero, etc.
In this paper, we propose a novel attention mechanism named Cog Attention [^1] The name is derived from the attention pattern, which resembles cogs. See Figure 10. that enables negative weights. Cog Attention exhibits superior properties from a mechanistic interpretability perspective and surpasses softmax attention in various applications without introducing any additional parameters. In Section 3, we provide mechanistic evidence for the expressiveness of Cog Attention:
(1) We identify attention heads that share the same working mechanism as exemplified above (Figures 1(b) and (d)), which shift the contextual process from the static OV matrix to dynamic QK inner products, with the OV matrix focusing more on refinement or modification. Irrelevant tokens are assigned negative weights for elimination, while other tokens are well preserved at the same time. This demonstrates Cog Attention’s enhanced flexibility and expressiveness compared to softmax attention.
(2) We demonstrate that models using Cog Attention exhibit improved robustness against representational collapse (Liu et al., 2020; Xie et al., 2023). Representational collapse refers to the phenomenon where representations become homogeneous in the later positions of a sequence within deep Transformer models. Barbero et al. (2024) contended that this issue arises because earlier tokens are “over-squashed” into later positions as the layer goes deeper. The negative weights in Cog Attention reduce the effective information paths from earlier tokens to later positions, thereby alleviating over-squashing and, consequently, mitigating representational collapse.
In Section 4, we develop Transformer-like models that use Cog Attention as attention modules and evaluate their performance across various tasks. Specifically, we train decoder-only language models for language modeling, and U-ViT diffusion models (Bao et al., 2023) for both unconditional and text-conditioned image generation. Our results show that across a wide range of tasks, these models equipped with Cog Attention achieve improved performance over the vanilla Transformer architecture using softmax attention.
2 Method
We begin by presenting the formulation of our proposed Cog Attention. Then, we discuss the design motivation and underlying principles.
2.1 Formulation
Let $\textbf{q},\textbf{k},\textbf{v}\in\mathbbm{R}^{n\times d}$ represent the query, key, and value vectors in an attention head. $n$ is the number of input tokens and $d$ is the hidden states dimension. The general attention computation can be expressed as follows:
$\textbf{p}_{i}$ $=\textbf{q}_{i}\textbf{k}^{\top},$ (1)
$\textbf{a}_{i}$ $=\phi(\textbf{p}_{i}),$
$\textbf{o}_{i}$ $=\sum^{i}_{j=0}\textbf{a}_{i,j}\textbf{v}_{j},$
where $\textbf{p}_{i}\in\mathbbm{R}^{1\times n}$ is the $i$-th row of the inner-product matrix. $\textbf{a}_{i}\in\mathbbm{R}^{1\times n}$ is the $i$-th row of the attention weights. $\textbf{o}_{i}\in\mathbbm{R}^{1\times d}$ is the weighted summation of vectors attended. [^2] Eq.1 is a causal attention formulation, as a token $i$ can only attend to preceding tokens $j\leq i$.
$\phi(\cdot)$ is the softmax function in a traditional attention module:
$\texttt{softmax}(\textbf{p}_{i})_{j}$ $=\frac{\texttt{exp}(\textbf{p}_{i,j}-m_{i})}{\sum^{i}_{k=0}\texttt{exp}(\textbf{p}_{i,k}-m_{i})},$ (2)
[…]
The subtraction of $m_{i}$ in the index aims to avoid numerical overflow.
In Cog Attention, we redefine $\phi(\cdot)$ as follows:
$\phi(\textbf{p}_{i})_{j}$ $=\frac{\texttt{SignExp}(\textbf{p}_{i,j})}{\sum^{i}_{k=0}\lvert\texttt{SignExp}(\textbf{p}_{i,k})\rvert},$ (3)
$\text{where}\ \texttt{SignExp}(\textbf{p}_{i,j})$ $=s_{i,j}\cdot\texttt{exp}(s_{i,j}\cdot\textbf{p}_{i,j}-m_{i}),$
$m_{i}=\texttt{max($\lvert\textbf{p}_{i}\rvert$)}\ \ \text{and}$ $\ \ s_{i,j}=\texttt{sign}(\textbf{p}_{i,j}).$
This formulation enables Cog Attention to produce negative attention weights. In the following subsection, we introduce and explain several design motivations and the underlying principles.
2.2 Design Principle
(1) The way to introduce negative weights. Although the inner product of query and key vectors naturally contains both positive and negative values, we apply an exponential function to this inner product and subsequently recover the sign of each term. This method is driven by our observation that an effective attention pattern for convergence must demonstrate sufficient kurtosis; i.e., it should be sparse and sharp enough. Without the exponential function, the attention pattern tends to be too flat, which can impede training convergence. We also tried using a cubic function as an alternative, which would eliminate the sign recovery process while still offering attention weights with adequate kurtosis. However, we ultimately chose the exponential function because of its convenience in gradient computation.
(2) The way to avoid numerical overflow. In Eq.3, we avoid numerical overflow in exponential functions by subtracting the maximum absolute value from the index. This approach differs from the softmax, which subtracts the maximum value, as seen in Eq.2. Our approach ensures that the maximum input to $\texttt{SignExp}(\cdot)$ remains 0, effectively avoiding overflow caused by a large $s_{i,j}\cdot\textbf{p}_{i,j}$, as illustrated in Figure 2.
Figure 2: The subtraction of the maximum absolute value in a row of query-key inner products avoids numerical overflow.
import torch
import torch.nn.functional as F
def Cog_Attention_naive(q, k, v, mask):
# q, k, and v’s shape: n x d
# mask shape: n x n, 0 for tokens to be masked, 1 for others
p = q @ k.transpose(1,0)
p = p * mask
p_sign = torch.sign(p)
max_p = torch.abs(p).max(-1, keepdim=True).values
e = p_sign * torch.exp(p_sign * p - max_p)
attn_w = e / torch.abs(e).sum(-1, keepdim=True)
return attn_w @ v
def Cog_Attention_fast(q, k, v, mask):
p = q @ k.transpose(1,0)
abs_p = torch.abs(p)
abs_p.masked_fill_(mask == 0, -float("inf"))
attn_w = torch.sign(p) * F.softmax(abs_p, dim=-1)
return attn_w @ v
Figure 3: A naive implementation of Cog Attention in Pytorch, alongside an equivalent yet faster implementation.
(3) The way to normalization. In softmax functions, the denominator in Eq.2 is the summation of the numerator, serving to normalize the outputs. This process ensures that the resulting attention weights sum to 1, with each term constrained within the range of $[0,1]$. Previous studies suggested that, for better convergence, the sum of attention weights in a row should remain stable, although not necessarily equal to a constant 1 (Wortsman et al., 2023). If not, training tricks are necessary to remain convergence and training stability (Ramapuram et al., 2024).
However, we challenge these findings. As shown in Eq. 3, normalizing by summing all outcomes of $\texttt{SignExp}(\cdot)$ may result in a zero denominator, causing NaN errors. To address this, we propose using the sum of the absolute values of the outcomes of $\texttt{SignExp}(\cdot)$ as the denominator. This adjustment leads to a non-constant summation across a row in the attention weight matrices. Nevertheless, our experiments demonstrate that Cog Attention maintains training stability and does not hinder convergence. This is because: (1) our method ensures that $\sum_{j=0}^{i}\lvert\phi(\textbf{p}_{i})_{j}\rvert$ remains constant at 1; and (2) based on the observation that the expectation of v is close to zero, adding or subtracting these value vectors—assumed to follow a multivariate Gaussian distribution—does not disrupt the norm expectation of the results, i.e., $\textbf{o}_{i}$ in Eq. 1.
We present a naive PyTorch implementation of Cog Attention, as shown in Figure 3, which closely follows the above description. Furthermore, we offer a more efficient equivalent implementation.
3 Enhanced Expressiveness of Cog Attention: A Mechanistic Interpretability Perspective
In this section, we provide mechanistic interpretability evidence to demonstrate that negative attention weights can enhance neural networks’ expressiveness.
3.1 Cog Attention Enhances the Flexibility of Attention Heads
Due to its unconstrained attention weights, Cog Attention enables processes such as deletion or copying, transitioning from a static OV matrix to dynamic query-key products. This capability stands as a key advantage of Cog Attention, facilitating concurrent processes within a single head, thereby enhancing the models’ flexibility and expressiveness.
We trained a Transformer language model with 141 million parameters and a Transformer-like language model using Cog Attention of the same size, respectively. Details regarding the model training can be found in Section 4. We studied attention heads’ working mechanisms on the indirect object identification (IOI) task (Wang et al., 2023), where the model is given with a context that includes the names of two people. For instance, given the input “Christopher and David had a lot of fun at school. David gave a ring to,” “Christopher” is the indirect object ($IO$), while “David” is the subject ($S$). The correct answer in IOI task is always the $IO$, which in this case is “Christopher.” There are 100 samples in the dataset.
To identify the most influential attention heads contributing to correct predictions in each model, we employed the path patching algorithm (Wang et al., 2023). We identified two significant heads: the 4th Cog Attention head in Layer 9 ($\text{CH}_{9.4}$) and the 11th softmax head in Layer 9 ($\text{SH}_{9.11}$). These heads accomplish the IOI task through a process of elimination. Figure 1(a) and (b) illustrate the attention weight patterns for these heads. Additionally, we computed the attention weights from the final token to both the $IO$ and $S$ tokens, plotted against the inner product of the heads’ outputs and $IO$ and $S$’s individual embedding, as shown in Figure 1(c) and (d).
The figures demonstrate that the two heads employ different mechanisms for the elimination process:
$\text{SH}_{9.11}$ assigns a large weight to the $S$ tokens, with its OV matrix (Elhage et al., 2021) generating a vector that opposes the representation of $S$’s embedding. Given that the function of the OV matrix is determined by the trained parameters, $\text{SH}_{9.11}$ also eliminates $IO$s due to its non-negative attention weights toward them. This is supported by Pearson correlation coefficient of $-0.21$ between the attention weights and the inner product, suggesting a weak to medium correlation. As a result, the non-negative nature of the softmax attention weights limits the performance on the IOI task.
In contrast, $\text{CH}_{9.4}$ assigns negative weights to $S$, effectively eliminating it while assigning minimal weights to $IO$s. Even if an $IO$ is assigned with a negative weight, we contend that the OV matrix in $\text{CH}_{9.4}$ is relieved from deletion to perform post-processing such as refinement or modification, and thus retains $IO$s. Evidence for this claim comes from the eigenvalue positivity [^3] Defined in the “Summarizing OV/QK Matrices” section in  (Elhage et al., 2021) for the OV matrices in two heads. The eigenvalue positivity, which indicates the OV matrix’s tendency toward copying (close to 1), deletion (close to -1), or abstract post-processing such as refinement or modification (close to 0), is 0.78 for $\text{CH}_{9.4}$ and -0.95 for $\text{SH}_{9.11}$. Additionally, the low Pearson correlation coefficient of $0.13$ between the attention weights and the inner product suggests little to no correlation between these two variables. As a result, $IO$s are not under “friendly fire” from $\text{CH}_{9.4}$.
3.2 Cog Attention is More Robust to Representational Collapse
Representational collapse (Liu et al., 2020; Xie et al., 2023) refers to that the representations at many positions become similar as the Transformer model goes deeper. Barbero et al. (2024) proposed an explanation for this issue, suggesting that earlier input tokens have more information paths to the final position compared to later tokens. As a result, the information from earlier tokens “over-squashed” into the representation at the final position, which is used for next-token prediction. This over-squashing causes representational collapse, leading Transformer models to struggle with distinguishing contexts that differ only slightly.
Figure 4: Two tasks for evaluating the extent of representational collapse in language models.
Figure 5: The $L_{+\infty}$ norms are computed for Transformer language models utilizing Cog Attention (subfigures (a) and (b)) and softmax attention subfigures (c) and (d)), respectively. The results are normalized by dividing them by the values obtained when $n=200$. Cog Attention makes language models more robust to representational collapse.
We demonstrate that Transformer models using Cog Attention are more robust to representational collapse than those using softmax attention. This improvement may be attributed to the negative weights reducing the number of effective information paths from earlier tokens to later positions, thereby mitigating the degree of over-squashing.
To assess the extent of representational collapse in a Transformer-based language model, we evaluate two tasks. In Task 1, “Finding a Zero,” the model processes two input sequences: the first consists of $n+1$ ones, and the second consists of $n$ ones with a leading zero. Following Barbero et al. (2024), we compute the $L_{+\infty}$ norm of the differences between the representations at the final positions of the two sequences. A larger $L_{+\infty}$ norm indicates that the model more effectively distinguishes between the two contexts, reflecting less representational collapse. Task 2, “Counting Ones,” follows a similar evaluation process.
We evaluate our trained 141M models using Cog Attention and softmax attention, respectively. The value of $n$ is set to 200, 400, 600, 800, 1000, and 2000, respectively.
Figure 5 illustrates the results. For each value of $n$ and each task, models employing Cog Attention exhibit significantly higher relative $L_{\infty}$ norms compared to those using softmax attention, highlighting Cog Attention’s robustness to representational collapse. This advantage of Cog Attention could be beneficial for tasks that involve long and complex contexts, such as retrieval-augmented generation (RAG) (Lewis et al., 2020).
Model ARC-E ARC-C PIQA SST-2 MNLI MRPC QQP RTE Avg.
Transformer 42.34 19.54 57.73 51.72 33.21 68.63 36.82 51.99 45.24
Cogformer 43.90 19.54 59.09 54.59 34.12 68.38 37.00 52.71 46.16
Table 1: Cogformer achieves a higher average accuracy on various tasks compared with Transformer models.
4 Experiments
We construct Transformer-like models using Cog Attention as the attention module, named Cogformer. We implement Cogformer in various tasks, including language modeling and image generation, to serve as a decoder-only model and a diffusion model, respectively. To showcase the effectiveness of Cog Attention  we evaluate and compare the performance of Cogformer with the Transformer across these diverse tasks.
4.1 Decoder-Only Models for Language Modeling
Architecture Details and Hyper-parameters. We train decoder-only Cogformer and Transformer-based language models on the RedPajama dataset (Computer, 2023). Apart from variations in the attention modules, the overall architecture and training hyper-parameters for Cogformer are identical to those of the vanilla Transformer. Importantly, we preserve the softmax attention in both the first and last layers of a deep Cogformer; we will elaborate on this design choice in Section 5.1. We use rotary position embedding (Su et al., 2023) and SwiGLU (Shazeer, 2020) activation in the feed-forward networks (FFN) following Llama (Touvron et al., 2023). RMSNorms are applied before both the attention and FFN modules. We adopt the Llama tokenizer, having a vocabulary of 32,000 tokens.
The model, with 141M parameters, consists of 12 layers, each with 12 attention heads. Its hidden state dimension is 768, while the intermediate dimension of the MLP layers is 3,072, with a context length of 2,048 tokens.
The models are trained on 100 billion tokens, with batch size of 131,072 tokens, an initial learning rate of 2e-4, and linear warmup for the first 2,000 steps followed by a cosine decay schedule to 4% of the peak learning rate. The AdamW optimizer (Loshchilov & Hutter, 2019) is configured with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a norm clipping value of 1, and a weight decay of 0.1. Our training is conducted on 8$\times$×A800-80G GPUs for approximately one week.
Test Tasks. We evaluate language models on a variety of widely used tasks that comprehensively assess their language modeling capabilities. These tasks include:
(1) ARC-easy (Clark et al., 2018), a multiple-choice question-answering task, and ARC-challenge, which presents greater challenges than ARC-easy.
(2) PIQA (Bisk et al., 2020), a task on physical commonsense reasoning.
(3) SST-2 (Socher et al., 2013), a binary sentiment classification task.
(4) MNLI (Williams et al., 2018), where the model predicts whether a given pair of sentences entail, contradict, or are unrelated to each other.
(5) MRPC (Dolan & Brockett, 2005), where the model assesses whether two sentences have equivalent semantics.
(6) QQP  [^4] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs, where the model determines whether two questions are semantically equivalent.
(7) RTE (GLUE (Wang et al., 2019) version), a two-class task for recognizing textual entailment.
All tasks are evaluated using accuracy as the metric, with all tests conducted in zero-shot, except for SST-2 which is one-shot tested. The evaluation code is based on LM Evaluation Harness (Gao et al., 2024).
Results. Table 1 presents the evaluation results. Cog Attention enhances the performance of Cogformer over the Transformer architecture across various tasks, highlighting its effectiveness in practical applications. Additionally, Figure 6 displays the training loss curves for the first 50,000 steps of both the Transformer and Cogformer architectures. The loss dynamics of the two models show no significant differences.
Figure 6: The convergence rates of Cogformer and the vanilla Transformer are nearly identical.
[…]
U-ViT-S/2 3.39 5.99
U-ViC-S/2 (Ours) 3.27 5.85
Table 2: Our U-ViC achieves better FID scores in unconditional image generation (CIFAR-10) and text-to-image generation (MS-COCO).
4.2 Diffusion Models for Image Generation
For our image generation tasks, we use U-ViT (Bao et al., 2023) as our baseline, which employs the Vision Transformer (Dosovitskiy et al., 2021) architecture as a diffusion model. We replace the softmax attention in U-ViT with Cog Attention (excluding the first and the last layers), resulting in our diffusion model, U-ViC. Both the U-ViC model and the U-ViT baseline are trained at a scale of 44 million parameters. The training is performed on the CIFAR-10 dataset (Krizhevsky, 2009) for unconditional image generation and the MS-COCO dataset (Lin et al., 2015) for text-conditioned image generation. For detailed information on the model architecture and training hyperparameters, please refer to Table 2 and Table 5 in (Bao et al., 2023). We evaluate performance using the FID score (Heusel et al., 2017), where lower values indicate better performance. Our experiments used 8 A800 GPUs, whereas Bao et al. (2023) used 4 × 2080 or 4 × A100 GPUs. Furthermore, we did not employ xFormer for acceleration. These differences in environments may contribute to the differences between the results reported in (Bao et al., 2023) and our reproduced ones.
Table 2 presents the results of our experiments. Our U-ViC model outperforms the U-ViT model across two datasets. Figure 7 illustrates generation samples.
Figure 7: Image generation samples. The figures are generated from the best checkpoint with lowest validation FID, and not cherry-picked.
5 Discussions
We discuss some important characteristics of Cog Attention, which might help the future research.
5.1 Convergence of Cogformer
As briefly mentioned before, in a deep Cogformer, we preserve the softmax attention in both the first and last layers. This aims to maintain the same convergence rate as a vanilla Transformer.
While we initially attempted to apply Cog Attention across all layers, we found that it resulted in slower convergence. Our observations of attention dynamics revealed that, at the beginning of training, the signs of attention weights do not accurately represent meaningful semantics due to the query-key inner product not being fully learned, leading to diverted optimization direction. The model appears to learn a shortcut at first, achieving lower training loss compared to the vanilla Transformer but then becoming stuck, allowing the vanilla Transformer to surpass it as training progresses. In contrast, softmax attention offers a near-uniform distribution that facilitates smoother initial training. Through empirical trials, we found that preserving softmax attention in the first layer of Cogformer effectively resolves the convergence issue.
Regarding the preservation of softmax attention in the last layer, we observed that a Cogformer with softmax attention only in the first layer resulted in last-layer attention weights’ sign highly consistent—either entirely positive or entirely negative. This suggests that the last layer may not require the expressiveness of negative weights. The training loss for these variants are shown in Table 3.
Model Training Step
500 1500 2500 3500 4500
Transformer 6.14 4.74 3.95 3.53 3.38
$\text{Cogformer}_{\text{all}}$ 6.07 4.73 4.03 3.60 3.45
$\text{Cogformer}_{\text{0}}$ 6.14 4.72 3.99 3.55 3.39
$\text{Cogformer}_{\text{0,last}}$ 6.15 4.73 4.02 3.53 3.38
Table 3: Training loss over the first 4500 steps for various models: $\text{Cogformer}_{\text{all}}$ (using Cog Attention in all layers), $\text{Cogformer}_{\text{0}}$ (applying softmax in the first layer), $\text{Cogformer}_{\text{0,last}}$ (applying softmax in both the first and last layers), and the vanilla Transformer (141M parameters).
5.2 Cog Attention Produces More Diverse Attention Patterns and Less Sink
Our analysis reveals that Cog Attention generates more diverse attention patterns compared to softmax attention. To investigate this, we examined the attention patterns across all heads in both Cogformer and Transformer, using the abstract section of “Attention Is All You Need” (Vaswani et al., 2017) as an input case. A comparison of Figures 9 and 10 shows that most heads in the vanilla Transformer exhibit sparse attention weights and a significant attention sink (Xiao et al., 2024), indicating that these heads are relatively inactive when processing the current input (Miller, 2023). In contrast, Cogformer displays more diverse attention patterns in its middle layers with reduced attention sink, suggesting that more heads are engaged in processing the input. We hypothesize that flexible attention patterns introduced by negative weights may lead to reduced parameter redundancy, but we currently lack direct evidence to quantify this reduction. Additionally, the diminished attention sink could potentially enhance extrapolation capabilities (Chen et al., 2024), KV cache compression (Liu et al., 2023b), high context-awareness task performance (Lin et al., 2024), and mitigate the lost-in-the-middle issue (Liu et al., 2023a). We will explore these questions in future research.
5.3 Time Cost Overhead
We compare the average time cost per training step on a single A800-80G GPU for our trained Transformer language model and Cogformer, using inputs of varying lengths. The results are shown in Figure 8. Due to the additional absolute operation and an extra multiplication, Cogformer incurs extra time overhead compared to the Transformer. We recognize this as a current limitation of our method and are actively exploring strategies to enhance its efficiency.
Figure 8: Comparison of time cost per training step for inputs of varying lengths between Transformer and Cogformer.
6 Related Works
Numerous efforts modified softmax function in the attention mechanism mainly for efficiency purposes, yet these attempts have not removed the constraints for non-negative weights. Some studies have proposed softmax alternatives such as ReLU (Shen et al., 2023; Wortsman et al., 2023), sigmoid (Ramapuram et al., 2024), cosine-based distance re-weighting (Qin et al., 2022), and learnable activations (Liu et al., 2024).
Recent studies showed the potential of an attention mechanism tailored to incorporate negative weights. Gong et al. (2024) demonstrate that removing the softmax function from the Mixture of Experts (MoE) router does not negatively impact performance. Ye et al. (2024) introduced differential attention, which calculates final attention weights by subtracting the outputs of two attention heads, thereby aiming to reduce noise. It might produce a small amount of negative attention weights as a by-product of the subtraction. It requires careful tuning of the scaling factor between the two heads and the inclusion of an additional normalization layer to ensure convergence. In this paper, we proposed Cog Attention, which introduces negative weights without requiring extra parameters or meticulous hyperparameter tuning. We demonstrated that Cog Attention increases the expressiveness of the attention mechanism.
7 Conclusions
In this paper, we introduced Cog Attention, a novel attention mechanism that incorporates negative weights. Our analysis, grounded in mechanistic interpretability, reveal that negative weights in attention computation enhance the expressiveness of neural networks and increase robustness against representational collapse. We subsequently implemented Cogformer, a new variant of the Transformer that integrates Cog Attention as its attention layer, to train language models and image generation diffusion models. The results indicate that Cog Attention can enhance Transformer performance across these tasks. We also discussed several properties of Cog Attention that could be beneficial for various applications, which we will explore in future works.
Acknowledgements
We appreciate the valuable discussions with Shen Nie regarding the U-ViT models, and to Yining Qian, Jia-Nan Li, Songhao Wu, and Shuqi Li for their helpful writing suggestions.